two machines that fill cereal boxes work at the same rate. after they work together for 7 h, one machine breaks down. the second machine requires 14 h more to finish filling the boxes. how long would it have taken one of the machines, working alone, to fill the boxes?
Answer by nerdybill(7384)

two machines that fill cereal boxes work at the same rate. after they work together for 7 h, one machine breaks down. the second machine requires 14 h more to finish filling the boxes. how long would it have taken one of the machines, working alone, to fill the boxes?
.
Let x = time (hours) it takes for one machine
then
7(1/x + 1/x) + 14(1/x) = 1
7(1 + 1) + 14(1) = x
7(2) + 14(1) = x
14 + 14 = x
28 hours = x